OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (magbak):

Hey @Hero @vinnv226 @satellite73 are you guys willing to help her in her question for me please. Thanks.

OpenStudy (jasmineflvs):

1) y = 5x + 3 y = 1/5x + 3 There is only one solution, because the lines intersect. 2) y = 5x + 3 y = 5x + 4 There are no solutions, because the lines are parallel. 3) y = 5x + 3 y = (10/2)x + 6/2 There are infinite solutions, because the lines are the same.

OpenStudy (jasmineflvs):

hope i helped

OpenStudy (anonymous):

Omg thank you SO much!!! @Jasmineflvs

OpenStudy (jasmineflvs):

no problem

OpenStudy (jasmineflvs):

any more questions i could answer

OpenStudy (anonymous):

Not at the moment but ill be sure to ask when i do! :)

OpenStudy (jasmineflvs):

okay

OpenStudy (jasmineflvs):

what grade are you in

OpenStudy (anonymous):

Ill be going into 8th but I'm doing online school to get ahead over the summer @Jasmineflvs

OpenStudy (jasmineflvs):

oh wow im behind in math im in 10th and im still in algebra 1...ughh